Cambridge A Level Mathematics 9709 — 2020 Oct/Nov Paper 5 · Variant 1
9709/51/O/N/20 · 6 questions · 50 marks · ≈56 min
The question paper and its mark scheme, free to read here and free to download. This is Cambridge’s own paper, exactly as it was sat.
Question paper12 pages












Mark scheme16 pages
Answers below. Sit the paper first if you are practising.
















Questions as text
Q2 · The probability that a student at a large music college plays in the band is 0.6
2 The probability that a student at a large music college plays in the band is 0.6. For a student who plays in the band, the probability that she also sings in the choir is 0.3. For a student who does not play in the band, the probability that she sings in the choir is x. The probability that a randomly chosen student from the college does not sing in the choir is 0.58. (a) Find the value of x. [3] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ Two students from the college are chosen at random. (b) Find the probability that both students play in the band and both sing in the choir. [2] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................
Mark scheme: 2(a) 0·6 × 0·7 + 0·4(1 – x) = 0·58 ≡ 0·42 + 0·4(1 – x) = 0·58 M1 Equation of form 0·6 × a + 0·4 × b = 0·58; a = 0·3, 0·7, b = x, (1 – x) B1 Single correct product seen, condone 0·42, in an equation of appropriate form x = 0·6 A1 Alternative method for question 2(a) 0·6 × 0·3 + 0·4x = 0·42 ≡ 0·18 + 0·4x = 0·42 M1 Equation of form 0·6 x a + 0·4 x b = 0·42; a = 0·3, 0·7, b = x, (1 – x) B1 Single correct product seen, condone 0·18, in an equation of appropriate form x = 0·6 A1 3 2(b) ( ) 2 0.6 0.3 × M1 (a × b)2, a = 0·6, 0·4 and b = 0·7, 0·3, x, (1–x) or 0·182, alone. 0.0324 A1 2 M1 pn, n = 6, 7 0 < p < 1
Q3 · Kayla is competing in a throwing event
3 Kayla is competing in a throwing event. A throw is counted as a success if the distance achieved is greater than 30 metres. The probability that Kayla will achieve a success on any throw is 0.25. (a) Find the probability that Kayla takes more than 6 throws to achieve a success. [2] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ (b) Find the probability that, for a random sample of 10 throws, Kayla achieves at least 3 successes. [3] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................
Mark scheme: 3(a) P(X > 6) = 6 0.75 0.178, 729 4096 A1 0·17797… 2 Question Answer Marks Guidance 3(b) 1 – P(0, 1, 2) = 1 – ( 10 0.75 + 10C1 1 9 0.25 0.75 + 10C2 2 8 0.25 0.75 ) M1 Binomial term of form 10Cx ( ) 10 1 , x x p p − − 0 < p < 1, any p, x ≠ 0, 10 1 – (0·0563135 + 0·1877117 + 0·2815676) A1 Correct unsimplified expression 0·474 A1 0·474 ⩽ p ⩽ 0·4744 3
Q4 · The random variable X takes each of the values 1, 2, 3, 4 with probability 1 Two…
4 The random variable X takes each of the values 1, 2, 3, 4 with probability 1 Two independent values 4. of X are chosen at random. If the two values of X are the same, the random variable Y takes that value. Otherwise, the value of Y is the larger value of X minus the smaller value of X. (a) Draw up the probability distribution table for Y. [4] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ (b) Find the probability that Y = 2 given that Y is even. [2] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................
Mark scheme: 4(a) y 1 2 3 4 prob 7 16 5 16 3 16 1 16 B1 1 2 3 4 1 1 1 2 3 2 1 2 1 2 3 2 1 3 1 4 3 2 1 4 Probability distribution table with correct scores with at least one probability, allow extra score values if probability of zero stated’ B1 One probability (linked with correct score) correct B1 2 more probs (linked with correct scores) correct B1 FT 4th prob correct, FT sum of 3 or 4 terms = 1 4 Question Answer Marks Guidance 4(b) P(2|even) = 5 16 6 16 M1 ( ) ( ) ( ) P 2 P 2 P 4 their their their + seen or correct outcome space. 5 6 or 0·833 A1 2
Q5 · The time in hours that Davin plays on his games machine each day is normally distributed…
5 The time in hours that Davin plays on his games machine each day is normally distributed with mean 3.5 and standard deviation 0.9. (a) Find the probability that on a randomly chosen day Davin plays on his games machine for more than 4.2 hours. [3] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ (b) On 90% of days Davin plays on his games machine for more than t hours. Find the value of t. [3] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ (c) Calculate an estimate for the number of days in a year (365 days) on which Davin plays on his games machine for between 2.8 and 4.2 hours. [3] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................
Mark scheme: 5(a) P(X > 4.2) = P( 4.2 3.5) 0.9 z − > = P( 0.7778) z > M1 Using ± standardisation formula, no 2 or σ σ , continuity correction 1 – 0·7818 M1 Appropriate area Φ , from standardisation formula P(z>…) in final solution 0·218 A1 3 5(b) 1.282 z = − B1 ±1.282 seen (critical value) 3.5 1.282 0.9 t − = − M1 An equation using ±standardisation formula with a z-value, condone 2 , σ σ and continuity correction 2.35 t = A1 AWRT, only dependent on M mark 3 Question Answer Marks Guidance 5(c) P(2.8 < X < 4.2) = 1 – 2 × their 5(a) ≡ 2(1 – their 5(a)) – 1 ≡ 2(0·5 – their 5(a)) = 0·5636 B1 FT FT from their 5(a) < 0.5 or correct Accept unevaluated probability OE Accept 0·564 Number of days = 365 × 0·5636 = 205·7 M1 365 × their p So, 205 (days) A1 FT Accept 205 or 206, not 205·0 or 206·0 no approximation/ rounding stated FT must be an integer value Alternative method for question 5(c) P 2.8 3.5 4.2 3.5 0.9 0.9 z − − < < ( ) ( ) Φ 0.7778 1 Φ0.7778 = − − = 0·7818 – (1 – 0·7818) = 0·5636 B1 0·5635 < p ⩽ 0·564 OE Number of days = 365 × 0·5636 = 205·7 M1 365 × their p So, 205 (days) A1 FT Accept 205 or 206, not 205·0 or 206·0 no approximation/ rounding stated FT must be an integer value 3
Q6 · The times, t minutes, taken by 150 students to complete a particular challenge are…
6 The times, t minutes, taken by 150 students to complete a particular challenge are summarised in the following cumulative frequency table. Time taken (t minutes) t ≤20 t ≤30 t ≤40 t ≤60 t ≤100 Cumulative frequency 12 48 106 134 150 (a) Draw a cumulative frequency graph to illustrate the data. [2] (b) 24% of the students take k minutes or longer to complete the challenge. Use your graph to estimate the value of k. [2] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ (c) Calculate estimates of the mean and the standard deviation of the time taken to complete the challenge. [6] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................
Mark scheme: 6(a) M1 At least 4 points plotted at upper end points, with both scales linear with at least 3 values indicated Correct cumulative frequency curve A1 All plotted correctly with curve drawn joined to (0, 0), axes labelled cumulative frequency, time, minutes 2 6(b) 150 × 0·76 = 114 M1 114 SOI, may be on graph k = 45 (mins) A1 FT Clear indication that their graph has been used, tolerance ±1mm 2 Question Answer Marks Guidance 6(c) Frequencies: 12 36 58 28 16 B1 Correct frequencies seen Mean = 10 12 25 36 35 58 50 28 80 16 150 × + × + × + × + × B1 At least 4 correct midpoints seen and used 120 900 2030 1400 1280 150 + + + + M1 Correct formula with their midpoints (not upper boundary, lower boundary, class width or frequency density). 38.2, 38 1 5 A1 Variance = 2 2 2 2 2 2 12 10 36 25 58 35 28 50 16 80 150 mean × + × + × + × + × − = 2 1200 22500 71050 70000 102400 150 mean + + + + − M1 Substitute their midpoints and frequencies (condone use of cumulative frequency) in correct variance formula, must have ‘– their mean2’ (Standard deviation = 321.76 ) = 17.9 A1 6
Q7 · Find the number of different ways in which the 10 letters of the word SHOPKEEPER can be…
7 (a) Find the number of different ways in which the 10 letters of the word SHOPKEEPER can be arranged so that all 3 Es are together. [2] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ (b) Find the number of different ways in which the 10 letters of the word SHOPKEEPER can be arranged so that the Ps are not next to each other. [4] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ (c) Find the probability that a randomly chosen arrangement of the 10 letters of the word SHOPKEEPER has an E at the beginning and an E at the end. [2] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ Four letters are selected from the 10 letters of the word SHOPKEEPER. (d) Find the number of different selections if the four letters include exactly one P. [3] ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................ ........................................................................................................................................................
Mark scheme: 7(a) 8! 2! M1 ( ) 8! 7! 8 ! ,where , , where 2 ! a k a k k × ≡ ∈ ∈ 20160 A1 2 Question Answer Marks Guidance 7(b) Total number of ways: ( ) 10! 302400 2!3! = (A) B1 Accept unsimplified With Ps together: 9! ( 3! = 60 480) (B) B1 Accept unsimplified With Ps not together: 302 400 – 60 480 M1 10! 9! m n − , m, n integers or (A) – (B) if clearly identified 241 920 A1 Alternative method for question 7(b) 8! 3! B1 k × 8! in numerator, k a positive integer, no ± B1 m × 3! in denominator, m a positive integer, no ± 9 8 2 × × M1 Their 8! 3!multiplied by 9C2 or 9P2 no additional terms 241 920 A1 Exact value, WWW 4 Question Answer Marks Guidance 7(c) Probability = Number of ways Es at beginning and end Total number of ways Probability = 8! 2! 10! 2! 3! × = 20160 302400 M1 8! ! 10! ! ! k k l 1 ⩽ k, l ∈ ℕ ⩽ 3, FT denominator from 7(b) or correct 1 15 , 0·0667 A1 Alternative method for question 7(c) Probability = 3 2 10 9 × M1 1 3,2 10 9 a a a − × = 1 15 , 0·0667 A1 Alternative method for question 7(c) Probability = 1 1 3! 10 9 × × M1 1 1 !, 3,2 10 9 m m × × = 1 15 , 0.0667 A1 2 Question Answer Marks Guidance 7(d) Scenarios: P E E E 5C0 = 1 P E E _ 5C1 = 5 P E _ _ 5C2 = 10 P _ _ _ 5C3 = 10 M1 5Cx seen alone, 1 ⩽ x ⩽ 4 M1 Summing the number of ways for 3 or 4 correct scenarios (can be unsimplified), no incorrect scenarios Total = 26 A1 3
What was in this paper
The subtopics covered by these 6 questions, and how many questions each got. Open one in a new tab to see every Cambridge question on it.
What you needed in this session
Cambridge’s own grade thresholds for 2020 Oct/Nov, Paper 5 · Variant 1. A higher threshold means an easier paper — the bar moves with how the cohort did.